Presentation is loading. Please wait.

Presentation is loading. Please wait.

MOM Report Paul Soler MICE Operations Manager University of Glasgow MICE Video Conference April 22, 2010.

Similar presentations

Presentation on theme: "MOM Report Paul Soler MICE Operations Manager University of Glasgow MICE Video Conference April 22, 2010."— Presentation transcript:

1 MOM Report Paul Soler MICE Operations Manager University of Glasgow MICE Video Conference April 22, 2010

2 MOM Report-Soler2 Decay Solenoid —Linde refrigerator developed fault: impurity levels rose due to a suspected blockage —Linde engineer Matthias Hoberich, visited RAL 23-26 March: changed charcoal in gas panel adsorber, heated it out, purged it and refilled compressor and no leak was found. —But when the compressor was isolated from cold box and ran system with cold trap in circuit the N 2 level rose from 50 ppm by 2-3 ppm/hour. —Since changing charcoal had no effect, it was suspected that the compressor shaft seal developed a leak and needed replacing —The collaboration meeting and Easter got in the way of a speedier resolution —Linde sent their engineer again between Wed-Fri 14-16 April). —HPC (local agents for Linde) received shaft seal Mon 12 April and sent engineer on Wed 14 April. —HPC tried to extract seal on Wed 14-15 April but it was lodged inside and would not move. —HPC had to dismantle compressor, take it back and have it removed on the bench. —They returned Tuesday 20 April with a new compressor since it is suspected that the fact that the seal was stuck on compressor shaft might have been cause of the problem.

3 April 22, 2010MOM Report-Soler3 Decay Solenoid Decay Solenoid (cont.) —Compresor (cont.): o New compressor was installed Tuesday and vacuum was reinstated – so far leak tests successful (vacuum ~2x10 -5 mbar) o Today and tomorrow: further purging and checking of impurity levels o When purging finished (hopefully tomorrow?), it will take ~10 days to cool down magnet – earliest will be 1 st week May. —Another problem with PTR90 Pirani/Penning vacuum gauges (giving false readings 4E-3 mbar) – 9 April o Examined with Leybold rep but they were not significantly contaminated o These were cleaned, they now work, but do not really understand why o Decision to purchase four more PTR90 heads for spares (£703 each) —Scroll pumps have never been serviced o DS south side pump struggling to pull vacuum below 3mbar – removed for servicing o DS north side pump can cope below 2E-3 mbar, so that pumping stack now in full use and pumping the DS! o Transfer line pump removed and replaced DS south side pump but also needs servicing and can only go down to 0.2 mbar

4 April 22, 2010MOM Report-Soler4 Target —Investigation of target forces —Emily Longi working to find if magnetic centre = geometric centre Magnetic Centre

5 April 22, 2010MOM Report-Soler5 Target Mechanical/Optical Measurements Bearing Tool Hall Probe A B Bore D C OpticalFaro A5.6 +/-0.1  AB=0.33 B5.89 +/-0.06 C6.57 +/-0.05  CD=0.94 D7.47 +/-0.05 Target —Mismatch between geometric and magnetic centre: ~0.5 mm —Is this the smoking for lateral forces in target shaft?

6 April 22, 2010MOM Report-Soler6 Target Forces —Simulations from Paul Smith on forces

7 April 22, 2010MOM Report-Soler7 Detectors —Luminosity Monitor o LM stand was bolted to ISIS floor during April shutdown o Dead time of LM: change LM gate from 10-40 ns and beamloss 0.5-2.0V) o Tried to carry it out on Friday 16 April but beam conditions were bad and there was a DAQ (human operational) error so no useful data. o Successful run on Tuesday 20 April (runs 1538-1555) – data is now being looked at —Tracker o Leak problem in cryostat due to cold weather over Christmas period o Leak was fixed last week (Alan Bross). o Sumitomo engineer replaced heads and now ready to cool down o Will debug DAQ before cooling down

8 April 22, 2010MOM Report-Soler8 Detectors and beam Detectors —CKOV,TOF0, TOF1, TOF2, KL: all powered up and ready to go o Request to have new TOF calibration with defocused beam to populate counters on the edge of acceptance o We will run ~10,000 actuations this weekend 23-26 April for this study (will the statistics be sufficient?) —Beam Profile Monitors: also ready to go Beam Magnets —Willie installed new pressure valves in D2, Q4-Q9 warm magnets. —Pressure in system reduced from ~7-8 bar to ~4 bar —Flow rate is now 60-67 l/min in quadrupoles Q4-Q9 and dipole D2 (used to be around 30-32 l/min). —Q1-Q3 and D1 still remain with flow rate ~30 l/min —Magnet polarity labelling scheme from Henry has been adopted Hall work: —PPS electrical system installation has started with DL staff —Nearly all planned jobs carried out and finished during shutdown (ie. cages now painted yellow, …)

9 April 22, 2010MOM Report-Soler9 Preparation for run Proposal for shifts during users run: —Inefficient to set up runs (~ 2 hours preparation) if work being carried out in hall —Can only work in evening (from 18:00 onwards) —Proposal to have a few evening shifts but to have bulk of running over weekends between 18:00 Friday to 7:00 Monday —We would need to have at least two shifts each day with two people on shift (for safety).

10 April 22, 2010MOM Report-Soler10 Run Plan Machine development period: 12-19 April —One shift (18:00-1:30) on Friday 16 April for Luminosity Monitor gate width studies: all hardware worked well but bad beam Programme ISIS run: 20 April – 28 May —Shift 20 April: rescheduled from Friday for LM gate studies o Run successful: took data 10, 20, 30, 40 ns gates for ~0.4 V, 1 V and 1.3 V – being analysed now —While decay solenoid not operational (20 April – first or second week May?) o Perform new TOF0,1,2 calibration with blown up defocused beam to try and cover all counters – 300 MeV pi+ beam with defocused Q7, Q8, Q9 by 20%, 20%, 40% (email from Marco) – 10,000 actuations scheduled 23-26 April o Proton absorber tests – move different thickness of absorber to compare particle rate vs absorber thickness – suggestion for weekend 30 April-3 May

11 April 22, 2010MOM Report-Soler11 Run Plan Programme ISIS run: 20 April – 28 May —While decay solenoid operational (First or second week May-28 May): o Perform target scan: delay target vs depth (ie acceleration) to try and flatten spill and reduce pileup and beam overlaps.- suggestion for weekend of 7-10 May (if DS ready! Is it worth while to do even if DS not ready?) o Work our way through Matrix of emittance (4 , 6 , 10  ) vs momentum (160, 200, 240 MeV/c) with all detectors on – 15-17 May, 22-24 May o Statistics for each point? ~10 5 muons? ISIS shutdown: 29 May-6 June Next machine development period: 7-21 June —Request further activation run: 16 hours with 4V beamloss? ISIS users run: 22 June – 12 August —Continue with Step I programme of emittance measurements vs momentum matrix. Machine Physics: 13-15 August Long shutdown: 16 August 2010-30 January 2011

Download ppt "MOM Report Paul Soler MICE Operations Manager University of Glasgow MICE Video Conference April 22, 2010."

Similar presentations

Ads by Google